A sample of 9 data yields the following results: sample mean is 4 and sample variance is 3. For unknown reasons, two of the sample values ​​are lost and when redoing the accounts with the remaining data, a new sample mean of 4 is obtained. and a new sample variance of 3/2.

a. Lost data has other values.
b. The missing data is 2 and 6.
c. The missing data is 0 and 8.
d. The missing data is 4 and 4.
A factory produces pistons whose diameters follow a normal distribution with a mean of 50 mm and a standard deviation of 0.01 mm. For a piston to serve, its diameter must be between 49.98 and 50.02 mm. If the diameter is less than 49.98 mm it is rejected; if it is greater than 50.02 mm, it is reprocessed once and the new diameter follows a normal distribution of an average of 49.99 mm, a standard deviation of 0.01 mm.

I. The probability that the piston must be reprocessed is 0.0228.

II. The probability that a reprocessed piston will be rejected is 0.61

a. Only I is correct

b. Only II is correct.

c. None is correct.

d. Both are correct.
Answer #1

Part~1:~Since~sample~variance~of~9~data=\frac{1}{8}\sum_{i=1}^9(x_i-\bar{x})^2=3\\ then~\sum_{i=1}^9(x_i-4)^2=8*3=24\\ Assume~x_8,~x_9~are~missing.~If~x_8=2,~x_9=6~then\\ \sum_{i=1}^7(x_i-4)^2=24-8=16~then~new~sample~variance=\frac{1}{6}\sum_{i=1}^7(x_i-4)^2\\ =16/6=8/3.\\\\ Hence~option~b~is~not~correct.\\\\ Again~assume~x_8=0,~x_9=8~then~\sum_{i=1}^7(x_i-4)^2=24-32<0~hence\\ Option~c.~is~incorrect.\\ assume~x_8=x_9=4~then~\sum_{i=1}^7(x_i-4)^2=24~then\\ new~sample~variance=24/6=4~hence\\ Option~d.~is~incorrect.\\ Hence~correct~option~a.

Part 2:

Probability that the piston must be reprocessed=P(X>50.02)=0.0228

R code: round(1-pnorm(50.02,50,0.01),4)

The probability that a reprocessed piston will be rejected=P(Y<49.98)=0.1587

R code: round(pnorm(49.98,49.99,0.01),4)

Option: a. Only I is correct